# Safe Learning for Contact-Rich Robot Tasks: A Survey from Classical Learning-Based Methods to Safe Foundation Models

### v1

@@ 3.2.1 Assembly and Insertion
tasks have been studied in the control literature for a long time both for their importance in industry and the challenges they offer [jiang2020state]. Robotic solutions to this problem can enable flexible assembly applications [67], and human-like activities such as inserting a key or a plug [179, 85].
The challenges include the precision requirement to achieve the tight fit, safety requirements to avoid damaging the equipment or getting stuck, and sensing difficulties as vision modality is less informative due to occlusions and the force-based nature of the task. Assembly problems may involve multi-step insertion actions to achieve a more complex plan [43, 10]. The classical peg-in-hole task focuses on a single insertion of a securely grasped peg into a single upright hole [levine2014learning]. The problem is gradually made more challenging through limited or noisy sensor data, leading to uncertainties over the peg and hole poses [104], varying peg shapes and sizes [87], or limited control options, such as rigid position-control [96].
@@ 3.2.2 Surface Interaction
tasks require a continuous purposeful interaction with a physical surface. This problem is important to achieve both autonomous manipulation tasks that are hazardous for humans such as sanding [89], and daily living activities in human spaces such as cleaning a table [35] or stirring a porridge [191]. These tasks usually require keeping a stable contact with the surface and applying consistent force while also following a desired trajectory. Maintaining a precise force and velocity profile enables the consistent quality of results. The problem is made more challenging by varying the surface curvature [2, 142] or tracking complex motion trajectories [107]. The robot may be equipped with a power tool such as a polisher that causes significant vibrations [425online].
The goal of the surface interaction tasks is to achieve a change on the surface. We observe these tasks under 3 categories: 1. Surface alteringtasks that change the surface geometry by removing material, such as grinding and sanding [89, 142]; 2. surface non-alteringtasks that do not cause a physical change in the surface, such as wiping and polishing [66, 56]; 3. particle interactiontasks that go beyond the surface and move particles, such as scooping a soft tissue [188] or stirring a granular medium [92, 191]. These tasks may require intricate 6-dof maneuvers and high force to move through or process the particles [12]. In case of deformable objects [zhu2022challenges], surface altering is easily achieved and revertible; in case of elastic objects such as a sponge [44], shape deformation is temporal and naturally reversed. Massaging is a special surface interaction task that is challenging due to involving a human, thus the safety concerns are more critical [46]. Interaction with deformable surfaces is an under-explored area, attracting more contributions.
We also identify the surface exploration tasks as a type of surface non-altering tasks that are less explored in the literature. In these works, the touch is fundamental to facilitate navigation and direct the agent towards the goal, such as in blind maze exploration [zhang2024srl, 375]. These tasks are similar to other surface interaction tasks as the interaction is usually non-prehensile and continuous, however, the goal of these tasks is not manipulating the surface, but navigating or following a trajectory [90].
@@ 3.2.3 Object Manipulation
The object manipulation tasks are commonly categorized as prehensile and non-prehensile. The former requires grasping an object, while the latter interacts with the object without holding it, such as pushing [94, 99], sliding [81] and pivoting [145]. Non-prehensile tasks can be done with a simpler manipulator that maintains a single contact with the object, however, they also have a limited control over the object motion. Consequently, the safety of these tasks depend on the motion of the object in challenging environmental conditions: [99] evaluate their method in cluttered environments where collisions are frequent and the object should be moved through narrow paths; the experiments of [94] involve pushing the object through a narrow gap where the object orientation becomes important, or through a bridge with the risk of the object falling. On the other hand, prehensile tasks require at least two contacts with the object to maintain a grasp. Simple grippers can tightly grasp an object, however they do not allow complex object manipulations, thus they are limited to pick-and-place or tool-use tasks. In contrast, bimanual manipulation is done through two high-DoF manipulators, and it can achieve complex interactions with the object such as lifting, pushing, flipping and rotating [22, 53]. Safe bimanual manipulation requires maintaining safe levels of contact forces without dropping the object, in addition to the workspace and motion limits [22]. Multi-fingered hands allow high manipulability of the object with increased control complexity. For a detailed review on dexterous hand manipulation we direct the reader to another survey [huang2025human].
The object manipulation tasks can be further categorized w.r.t. the nature of the object, such as constrained, articulated, and deformable. Constrained object manipulation tasks, such as opening a door or a drawer, impose strict kinematic constraints on the object motion [86]. These constraints may lead to high levels of force, if not handled with compliant and adaptive behavior [36]. Objects like a door, drawer or lever are also termed articulated objects [7]. Disassembly tasks (Sec. 3.2.1) are challenging examples of constrained object manipulation. Deformable object manipulation is an under-explored research problem that is challenging due to the sensing and modeling complexities [zhu2022challenges].
@@ 3.2.4 Physical HRI
tasks involve physical contact with humans [52], unlike the human-robot coexistence tasks, where avoiding the human is the main goal of the safety method [57]. pHRI tasks are challenging because of the increased criticality of human safety and the added uncertainty due to the human actor. Thus, the methods should estimate the human behavior and accommodate the uncertainty [ajoudani2018progress]. In traditional pHRI, the robot assumes the passive role, and follows the human intention through compliance control. A special case of this is a robot interacting with both the human and the environment, as in [17], where the human guides the robot while it interacts with a surface. In [59], the robot can switch between human-guided and autonomous in-contact manipulation modes to adapt its impedance characteristics.
The safety-critical contact-rich interactions between an active robot and human fit better to our scope. In prominent scenarios, the robot touches the human while the human takes a passive role, such as dressing [9] and bathing assistance [5], ultrasound scanning [14], and massaging [19, 46]. When the robot is physically coupled with the human, such as the rehabilitation robotics, the controller should estimate the human impedance parameters to ensure stability [24].
Crucially, the notion of safety expands with the inclusion of a human. For example, [5] limit the vision system to a LIDAR during bathing assistance for ethical considerations. Human perceived safety also becomes important for the trustworthy and socially acceptable interactions [bartneck2009measurement].
@@ 3.3 Sensing and Policy Modalities
Proprioceptive and kinematic state only Contact reasoning from joint configurations and end effector pose; conservative motion generation [21, 27, 34, 42, 45, 81, 90, 96, 99, 101, 103]
Force/torque Real-time enforcement of force and impedance limits, slip detection, compliant correction of model error [1, 2, 3, 4, 6, 7, 9, 10, 11, 12, 14, 15, 16, 17, 18, 19, 20, 23, 25, 26, 29, 30, 32, 35, 36, 37, 39, 40, 41, 44, 46, 49, 55, 56, 58, 59, 60, 61, 63, 64, 66, 67, 68, 69, 70, 72, 73, 74, 75, 76, 77, 78, 79, 80, 88, 92, 94, 95, 97, 98, 100, 104]
Vision Contact evaluation, localizing task features, and supplying global context that force sensors cannot provide [1, 2, 3, 5, 6, 9, 12, 13, 15, 18, 20, 22, 28, 33, 36, 47, 48, 57, 61, 65, 67, 68, 85, 87, 89, 91, 94, 95, 98, 100, 107]
Natural-language instruction Large multimodal models interpret a human’s insturction and then check internal low-level safety [8, 15, 33, 51]
Tactile High-resolution contact localization, pressure distribution monitoring, and incipient slip detection [43, 71, 86, 106]
Multimodal fusion Enables high precision under occlusion [1, 2, 3, 6, 9, 12, 15, 18, 20, 33, 36, 61, 67, 68, 71, 94, 95, 98, 100, 106]
@@ 3.3.1 Pose and Proprioceptive State
Pose-dominant observations are also used when exteroceptive information is represented directly in configuration space rather than as raw pixels. Benchmark interaction tasks encode the positions of pucks, tools or human partners as Cartesian poses, combined with the robot’s own pose and velocity, so that learning algorithms can enforce distance-based constraints, safety fields around humans, or recovery zones in a low-dimensional geometric space [91, 48, zhang2025bresa]. Even in visuomotor settings where images are available, the safety shaping is often defined over the proprioceptive and pose coordinates through joint limits, collision distances or certified invariant sets, while the visual input mainly serves to infer these geometric quantities or to disambiguate task context [13, 47, 95, 99, 107]. Compared with force-centric designs, such pose-based formulations trade direct measurement of contact forces for simpler, more structured safety reasoning in configuration space, and are particularly attractive when high-quality kinematic models and planners are available.
@@ 3.3.2 Force and Torque Sensing
For many contact-rich tasks, especially those involving insertion, surface following, or physical human–robot interaction, F/T measurements at the wrist or fingers act as the primary safety interface. Policies typically observe these signals together with joint and end-effector state and use them to cap contact loads, detect slip or jamming, and adapt impedance so that tracking performance is achieved without violating force or pressure limits [1, 2, 4, 23]. In contrast to purely kinematic designs, such formulations expose safety-relevant quantities directly in the observation, which allows controllers to react immediately to unsafe contact events in assembly, surface interaction, and terrain-contact scenarios [19, 32, 39, 61].
@@
F/T signals are also combined with explicit safety margins, constraints or stability certificates derived from control theory. In this line of work, the observation space does not stop at raw wrench measurements and state. Controllers additionally receive features derived from Lyapunov functions, control barrier functions, passivity errors or other safety indicators that are computed using the measured forces and state trajectories [14, 16, 29, 59]. Learning algorithms then operate directly on these structured safety variables, for example by penalizing trajectories that approach constraint boundaries, adjusting admittance or impedance gains when passivity residuals grow, or biasing exploration away from regions with high estimated risk [39, 44, 57, 80]. This idea extends naturally to uncertainty-aware formulations, where system state, force/torque and uncertainty estimates are combined, and the policy is trained to remain robustly within contact-safe sets despite modelling errors and partial observability [60, 92].
In physical human–robot interaction and teleoperation, F/T channels additionally encode human intent and comfort. Many frameworks treat human-applied forces at the end-effector, interaction torques, or teleoperated wrenches as part of the policy input together with robot joint or end-effector state and task phase [11, 17, 24, 58]. These signals are interpreted either as demonstrations of desired compliant behaviour or as instantaneous preferences over interaction forces, and are used to adapt impedance online, to learn tele-impedance mappings, or to construct task-oriented safety fields around the human partner [15, 52, 69, 76]. Safety here is twofold: the robot must both respect human-specified force envelopes and ensure that the resulting closed-loop behaviour remains stable under changing human dynamics.
@@ 3.3.4 Language Instructions
Natural language is increasingly used as a high-level input channel through which humans specify tasks, constraints and preferences, while lower-level controllers remain responsible for enforcing physical safety during execution. Instead of hard-coding task objectives in cost functions or reward weights, these approaches let users describe what the robot should do, and often how cautiously it should behave, in free-form  [kyaw_text_2025], which is then grounded into symbolic plans, motion goals or policy parameters [8, 33, 51]. From the perspective of safety inputs, language thus complements geometric and haptic signals by exposing human intent and safety requirements explicitly at the policy interface.
One line of work couples language with visual and proprioceptive observations in large VLA architectures. In such systems, camera images and robot state are processed jointly with a natural-language instruction to produce either low-level actions or intermediate code that can be executed by downstream controllers [8, 33]. Safety is then imposed at two levels: the model is aligned to avoid generating obviously harmful plans, and the generated code or action sequences are checked against constraint sets or safety filters before deployment [8, 33]. Compared to purely numeric task encodings, this design allows the same sensorimotor stack to be reused across tasks, while safety-relevant differences are captured in the language input rather than in separate controllers.
@@
Across these works, language does not replace low-level safety mechanisms such as force limits, control barrier functions or conservative exploration. Rather, it provides a flexible interface for specifying and updating what safe means in a given context. High-level instructions determine which regions of the state and action spaces are desirable or forbidden, while the underlying visuomotor and force-aware controllers ensure that, once a plan is chosen, contact forces, distances and stability constraints remain within acceptable bounds [8, 33, 51, 54]. As language models and alignment techniques improve, this combination of linguistic intent with structured safety inputs offers a promising route to robots that can be both responsive to human guidance and reliably safe in contact-rich environments.

### v2

@@ 3.2.1 Assembly and Insertion
Tasks have been studied in the control literature for a long time both for their importance in industry and the challenges they offer [126]. Robotic solutions to this problem can enable flexible assembly applications [162], and human-like activities such as inserting a key or a plug [196, 280].
The challenges include the precision requirement to achieve the tight fit, safety requirements to avoid damaging the equipment or getting stuck, and sensing difficulties as vision modality is less informative due to occlusions and the force-based nature of the task. Assembly problems may involve multi-step insertion actions to achieve a more complex plan [172, 226]. The classical peg-in-hole task focuses on a single insertion of a securely grasped peg into a single upright hole [157]. The problem is gradually made more challenging through limited or noisy sensor data, leading to uncertainties over the peg and hole poses [23], varying peg shapes and sizes [281], or limited control options, such as rigid position-control [24].
@@ 3.2.2 Surface Interaction
tasks require a continuous purposeful interaction with a physical surface. This problem is important to achieve both autonomous manipulation tasks that are hazardous for humans such as sanding [117], and daily living activities in human spaces such as cleaning a table [60] or stirring a porridge [208]. These tasks usually require keeping a stable contact with the surface and applying consistent force while also following a desired trajectory. Maintaining a precise force and velocity profile enables the consistent quality of results. The problem is made more challenging by varying the surface curvature [167, 391] or tracking complex motion trajectories [205]. The robot may be equipped with a power tool such as a polisher that causes significant vibrations [274].
The goal of the surface interaction tasks is to achieve a change on the surface. We observe these tasks under 3 categories: 1. Surface alteringtasks that change the surface geometry by removing material, such as grinding and sanding [117, 391]; 2. surface non-alteringtasks that do not cause a physical change in the surface, such as wiping and polishing [399, 359]; 3. particle interactiontasks that go beyond the surface and move particles, such as scooping a soft tissue [286] or stirring a granular medium [146, 208]. These tasks may require intricate 6-dof maneuvers and high force to move through or process the particles [3]. In case of deformable objects [398], surface altering is easily achieved and revertible; in case of elastic objects such as a sponge [323], shape deformation is temporal and naturally reversed. Massaging is a special surface interaction task that is challenging due to involving a human, thus the safety concerns are more critical [170]. Interaction with deformable surfaces is an under-explored area, attracting more contributions.
We also identify the surface exploration tasks as a type of surface non-altering tasks that are less explored in the literature. In these works, the touch is fundamental to facilitate navigation and direct the agent towards the goal, such as in blind maze exploration [374, 112]. These tasks are similar to other surface interaction tasks as the interaction is usually non-prehensile and continuous, however, the goal of these tasks is not manipulating the surface, but navigating or following a trajectory [240].
@@ 3.2.3 Object Manipulation
The object manipulation tasks are commonly categorized as prehensile and non-prehensile. The former requires grasping an object, while the latter interacts with the object without holding it, such as pushing [331, 347], sliding [137] and pivoting [382]. Non-prehensile tasks can be done with a simpler manipulator that maintains a single contact with the object, however, they also have a limited control over the object motion. Consequently, the safety of these tasks depend on the motion of the object in challenging environmental conditions: [347] evaluate their method in cluttered environments where collisions are frequent and the object should be moved through narrow paths; the experiments of [331] involve pushing the object through a narrow gap where the object orientation becomes important, or through a bridge with the risk of the object falling. On the other hand, prehensile tasks require at least two contacts with the object to maintain a grasp. Simple grippers can tightly grasp an object, however they do not allow complex object manipulations, thus they are limited to pick-and-place or tool-use tasks. In contrast, bimanual manipulation is done through two high-DoF manipulators, and it can achieve complex interactions with the object such as lifting, pushing, flipping and rotating [138, 64]. Safe bimanual manipulation requires maintaining safe levels of contact forces without dropping the object, in addition to the workspace and motion limits [138]. Multi-fingered hands allow high manipulability of the object with increased control complexity. For a detailed review on dexterous hand manipulation we direct the reader to another survey [115].
The object manipulation tasks can be further categorized w.r.t. the nature of the object, such as constrained, articulated, and deformable. Constrained object manipulation tasks, such as opening a door or a drawer, impose strict kinematic constraints on the object motion [311]. These constraints may lead to high levels of force, if not handled with compliant and adaptive behavior [127]. Objects like a door, drawer or lever are also termed articulated objects [1]. Disassembly tasks (Sec. 3.2.1) are challenging examples of constrained object manipulation. Deformable object manipulation is an under-explored research problem that is challenging due to the sensing and modeling complexities [398].
@@ 3.2.4 Physical HRI
tasks involve physical contact with humans [73], unlike the human-robot coexistence tasks, where avoiding the human is the main goal of the safety method [369]. pHRI tasks are challenging because of the increased criticality of human safety and the added uncertainty due to the human actor. Thus, the methods should estimate the human behavior and accommodate the uncertainty [6]. In traditional pHRI, the robot assumes the passive role, and follows the human intention through compliance control. A special case of this is a robot interacting with both the human and the environment, as in [341], where the human guides the robot while it interacts with a surface. In [236], the robot can switch between human-guided and autonomous in-contact manipulation modes to adapt its impedance characteristics.
The safety-critical contact-rich interactions between an active robot and human fit better to our scope. In prominent scenarios, the robot touches the human while the human takes a passive role, such as eating [45], dressing [290] and bathing assistance [345], ultrasound scanning [313], and massaging [339, 170]. When the robot is physically coupled with the human, such as the rehabilitation robotics, the controller should estimate the human impedance parameters to ensure stability [261].
Crucially, the notion of safety expands with the inclusion of a human. For example, [345] limit the vision system to a LIDAR during bathing assistance for ethical considerations. Human perceived safety also becomes important for the trustworthy and socially acceptable interactions [19].
@@ 3.3 Sensing And Policy Modalities
Proprioceptive and kinematic state only Contact reasoning from joint configurations and end effector pose; conservative motion generation Mitsioni et al. [217], Wang et al. [325], Zhu et al. [397], Chen et al. [46], Günster et al. [94], Kim et al. [137], Peng et al. [240], Beltran-Hernandez et al. [24], Yamada et al. [347], Soni et al. [275], Thananjeyan et al. [302]
Force/torque Real-time enforcement of force and impedance limits, slip detection, compliant correction of model error Cramer et al. [56], Li et al. [167], Zang et al. [365], Okada et al. [232], Hou et al. [106], Abbatematteo et al. [1], Sun et al. [290], Noseworthy et al. [226], Padalkar et al. [234], Aburub et al. [3], Vinter-Hviid et al. [313], Zhu et al. [396], Li et al. [165], Xing et al. [341], Liu et al. [181], Xiao et al. [339], Qu et al. [242], Yang et al. [350, 351], Yan et al. [348], Khader et al. [132], Hou et al. [108], Li et al. [168], Dalle Vedove et al. [60], Kang et al. [127], Zhang et al. [381], Jadav and Palanthandalam-Madapusi [121], Anand et al. [11], Ji et al. [124], Wang et al. [323], Li et al. [170], Gai et al. [79], Chen et al. [47], Yu et al. [359], Giammarino et al. [83], Pan et al. [236], Yao et al. [354], Ding et al. [65], Xiang et al. [338], Wu et al. [334], Zhu et al. [399], Li et al. [162], Wang et al. [326], Anand et al. [9], Xu et al. [343], Li et al. [166], Zhang et al. [378], Liu et al. [175], Karacan et al. [131], Roveda et al. [249], Beltran-Hernandez [22], Feng et al. [74], Li et al. [163], Wang et al. [316], Zhao et al. [386], Kuo et al. [146], Wirnshofer et al. [331], Hou et al. [107], Bogdanovic et al. [32], Zhao et al. [385], Bahl et al. [14], Beltran-Hernandez et al. [23]
Vision Contact evaluation, localizing task features, and supplying global context that force sensors cannot provide Cramer et al. [56], Li et al. [167], Zang et al. [365], Xu et al. [345], Hou et al. [106], Sun et al. [290], Aburub et al. [3], Lancaster et al. [151], Zhu et al. [396], Liu et al. [181], Qu et al. [242], Kim et al. [138], Li and Xu [169], Zhang et al. [367], Kang et al. [127], Mayr [210], Chuck et al. [54], Zhang et al. [369], Ding et al. [65], Zhan et al. [366], Li et al. [162], Wang et al. [326], Spector and Di Castro [280], Spector and Zacksenhouse [281], Huo et al. [117], Thananjeyan et al. [301], Wirnshofer et al. [331], Hou et al. [107], Zhao et al. [385], Bahl et al. [14], Martín-Martín et al. [205]
Natural-language instruction Large multimodal models interpret a human’s insturction and then check internal low-level safety Ni et al. [225], Zhu et al. [396], Zhang et al. [367], Thumm et al. [303]
Tactile High-resolution contact localization, pressure distribution monitoring, and incipient slip detection Lin et al. [172], Zhang et al. [379], Ulmer et al. [311], Luu et al. [197], Huang et al. [114]
Multimodal fusion Enables high precision under occlusion Cramer et al. [56], Li et al. [167], Zang et al. [365], Hou et al. [106], Sun et al. [290], Aburub et al. [3], Zhu et al. [396], Liu et al. [181], Qu et al. [242], Zhang et al. [367], Kang et al. [127], Ding et al. [65], Li et al. [162], Wang et al. [326], Zhang et al. [379], Wirnshofer et al. [331], Hou et al. [107], Zhao et al. [385], Bahl et al. [14], Huang et al. [114]
@@ 3.3.1 Pose and Proprioceptive State
Pose-dominant observations are also used when exteroceptive information is represented directly in configuration space rather than as raw pixels. Benchmark interaction tasks encode the positions of pucks, tools or human partners as Cartesian poses, combined with the robot’s own pose and velocity, so that learning algorithms can enforce distance-based constraints, safety fields around humans, or recovery zones in a low-dimensional geometric space [301, 54, 372]. Even in visuomotor settings where images are available, the safety shaping is often defined over the proprioceptive and pose coordinates through joint limits, collision distances or certified invariant sets, while the visual input mainly serves to infer these geometric quantities or to disambiguate task context [151, 210, 107, 347, 205]. Compared with force-centric designs, such pose-based formulations trade direct measurement of contact forces for simpler, more structured safety reasoning in configuration space, and are particularly attractive when high-quality kinematic models and planners are available.
@@ 3.3.2 Force and Torque Sensing
For many contact-rich tasks, especially those involving insertion, surface following, or physical human–robot interaction, F/T measurements at the wrist or fingers act as the primary safety interface. Policies typically observe these signals together with joint and end-effector state and use them to cap contact loads, detect slip or jamming, and adapt impedance so that tracking performance is achieved without violating force or pressure limits [374, 56, 167, 232, 350]. In contrast to purely kinematic designs, such formulations expose safety-relevant quantities directly in the observation, which allows controllers to react immediately to unsafe contact events in assembly, surface interaction, and terrain-contact scenarios [339, 168, 121, 65].
@@
F/T signals are also combined with explicit safety margins, constraints or stability certificates derived from control theory. In this line of work, the observation space does not stop at raw wrench measurements and state. Controllers additionally receive features derived from Lyapunov functions, control barrier functions, passivity errors or other safety indicators that are computed using the measured forces and state trajectories [313, 165, 132, 236]. Learning algorithms then operate directly on these structured safety variables, for example by penalizing trajectories that approach constraint boundaries, adjusting admittance or impedance gains when passivity residuals grow [373, 364], or biasing exploration away from regions with high estimated risk [121, 323, 369, 316]. This idea extends naturally to uncertainty-aware formulations, where system state, force/torque and uncertainty estimates are combined, and the policy is trained to remain robustly within contact-safe sets despite modelling errors and partial observability [354, 146].
In physical human–robot interaction and teleoperation, F/T channels additionally encode human intent and comfort. Many frameworks treat human-applied forces at the end-effector, interaction torques, or teleoperated wrenches as part of the policy input together with robot joint or end-effector state and task phase [234, 341, 261, 83]. These signals are interpreted either as demonstrations of desired compliant behaviour or as real-time preferences over interaction forces, and are used to adapt impedance online, to learn tele-impedance mappings, or to construct task-oriented safety fields around the human partner [396, 73, 9, 249]. Safety here is twofold: the robot must both respect human-specified force envelopes and ensure that the resulting closed-loop behaviour remains stable under changing human dynamics.
@@ 3.3.4 Language Instructions
Natural language is increasingly used as a high-level input channel through which humans specify tasks, constraints and preferences, while lower-level controllers remain responsible for enforcing physical safety during execution. Instead of hard-coding task objectives in cost functions or reward weights, these approaches let users describe what the robot should do, and often how cautiously it should behave, in free-form  [148], which is then grounded into symbolic plans, motion goals or policy parameters [225, 367, 303]. From the perspective of safety inputs, language thus complements geometric and haptic signals by exposing human intent and safety requirements explicitly at the policy interface.
One line of work couples language with visual and proprioceptive observations in large vision–language–action architectures. In such systems, camera images and robot state are processed jointly with a natural-language instruction to produce either low-level actions or intermediate code that can be executed by downstream controllers [225, 367]. Safety is then imposed at two levels: the model is aligned to avoid generating obviously harmful plans, and the generated code or action sequences are checked against constraint sets or safety filters before deployment [225, 367]. Compared to purely numeric task encodings, this design allows the same sensorimotor stack to be reused across tasks, while safety-relevant differences are captured in the language input rather than in separate controllers.
@@
Across these works, language does not replace low-level safety mechanisms such as force limits, control barrier functions or conservative exploration. Rather, it provides a flexible interface for specifying and updating what safe means in a given context [367]. High-level instructions determine which regions of the state and action spaces are desirable or forbidden, while the underlying visuomotor and force-aware controllers ensure that, once a plan is chosen, contact forces, distances and stability constraints remain within acceptable bounds [225, 367, 303, 89]. As language models and alignment techniques improve, this combination of linguistic intent with structured safety inputs offers a promising route to robots that can be both responsive to human guidance and reliably safe in contact-rich environments.